OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 5x + 3y = 15? giving out medals

OpenStudy (anonymous):

5x-(3)(0)=-15 5x=-15 (5x)(1/5)=(-15)(1/5) x = -3 so the x intercept is (-3,0)